# EmoVoice: LLM-based Emotional Text-To-Speech Model with Freestyle Text Prompting

### v3

@@ 6.1.1. Objective evaluation results
The experimental results in Table 2 demonstrate that our proposed model, EmoVoice, achieves competitive performance across all evaluated metrics. Compared with existing emotional speech synthesis models, EmoVoice(1.5B) achieves the most balanced overall performance, with highest emotion similarity of 0.9118, and highest recall rate of 0.424, while maintaining low WER of 2.62 and high UTMOS score of 4.345. It rivals the powerful GPT-4o-mini-tts in content consistency and emotional expressiveness, while surpassing it in terms of speech quality. These results indicate that EmoVoice not only preserves linguistic accuracy and delivers high perceptual quality, but also excels in producing emotionally rich speech that faithfully represents the intended emotional directives. GPT-4o-audio is used to generate the test set and serves as an upper bound reference rather than a comparable model.
@@ 6.1.2. Subjective evaluation results
Table 3. Comparison of Mean Opinion Score (MOS) for emotional expressiveness and instruction-following capabilities of different emotion-controllable TTS models, evaluated by human raters.
@@
Given the inherently subjective and abstract nature of emotional expressiveness, objective metrics can only provide limited insight into model performance. Therefore, we conduct a subjective assessment to better align evaluations with human perception. Specifically, we randomly sample 10 utterances form each of six emotion categories—angry, happy, sad, surprised, disgusted, and fearful—from the test set, totaling 60 audio samples. These samples are evaluated by 30 distinct participants, each of whom rates 10 random groups of speech produced by six representative models. The evaluation focuses on two key aspects: the overall emotional expressiveness of the synthesized speech and its adherence to the given emotional instructions. Ratings are scored on a scale of 1 to 5, with increments of 0.5.
The results presented in Table 3 reveal that the Mean Opinion Score (MOS) trends align closely with the objective emotion metrics, validating the consistency of our findings. Notably, the GPT-4o series models exhibit the strongest performance, with GPT-4o-audio achieving the highest MOS of 4.35, reflecting its ability to produce highly compelling and emotionally rich speech. Our EmoVoice models demonstrate competitive performance, surpassing both CosyVoice2 and PromptTTS by a significant margin. Specifically, the EmoVoice(1.5B) achieves a MOS of 3.507, closely approaching the performance of GPT-4o-mini-tts of 3.598 and significantly outperforming CosyVoice2 of 2.138 and PromptTTS of 1.415, reaffirming the effectiveness of our approach in producing emotionally expressive and instruction-aligned speech.
@@ 6.1.3. Chinese model results
Table 4. Performance comparison of EmoVoice and other emotion-controllable TTS models in terms of WER, emotion similarity, emotion classification recall rate, and UTMOS on Secap test set.
@@ 6.2.1. Different output-side structures
Table 5. Performance comparison of EmoVoice variants with different output structures on EmoVoice-DB test set. Both Fine-tuned Emotional TTS in the second stage and Pre-trained TTS in the first stage are evaluated by WER, emotion similarity, emotion classification recall rate, and UTMOS. WER of the Pre-trained TTS is also computed on test-en of Seed-TTS (Anastassiou et al., 2024).
@@
EmoVoice-PP Parallel Phoneme and Audio Tokens 3.06 0.9115 0.379 4.35 3.94/3.11 0.8704 0.298 4.40
@@
We explore optimizing the output-side structure of the EmoVoice model, comparing six different architectures, including standard single-output stream and parallel output streams, as shown in Figure 2. We investigate utilizing text and phonemes as auxiliary information to guide the generation of audio tokens. The baseline EmoVoice model outputs a pure audio token sequence. The EmoVoice-I model produces an interleaved sequence of text and audio tokens, mixed in a predefined ratio of 12:36, a structure commonly employed in streaming TTS models (Du et al., 2024b; Yang et al., 2024b). Inspired by CoT and CoM techniques, we experiment with first predicting text or phoneme tokens as an auxiliary intermediate thought process before generating the target audio token within a single output stream. These models are denoted as EmoVoice-ST(Serial Text) and EmoVoice-SP(Serial Phoneme), respectively. Besides, we explore using an additional prediction head in the output structure to produce text or phoneme tokens simultaneously. EmoVoice-PT(Parallel Text) and EmoVoice-PP(Parallel Phoneme) represent models that parallelly predict audio tokens alongside text tokens and phoneme tokens, respectively.
@@ 6.2.3. Scaling LLM size
Given the well-established scaling law in LLMs, we explore how scaling the parameter size of the LLM backbone affects the model’s performance on the emotional TTS task. Specifically, we compare the performance of EmoVoice with two different sizes of 0.5B and 1.5B parameters, using the qwen2.5-0.5B and qwen2.5-1.5B models respectively.
As shown in Tabel 7, in the pre-training phase, we observe an obvious reduction of WER when scaling up the model. The 1.5B model achieves a WER of 3.83% with a relative reduction of 19.03% compared to the 4.73% WER of the 0.5B model, indicating enhanced content consistency. During the emotion fine-tuning phase, the recall rate rises from 0.395 to 0.424 and the WER reduces from 2.73% to 2.62%. Emotion similarity and UTMOS remain largely consistent. Additionally, subjective evaluation results in Table 3 show the MOS score improves from 3.163 to 3.507. Overall, scaling up the LLM effectively enhances emotional expressiveness and content consistency. However, it incurs substantial additional computational costs and training time, with improvements constrained by the limited emotion dataset size.
@@ 6.2.4. LLM Initialization
While several TTS models adopt LLMs as their backbone and utilize their parameters for initialization, the specific impact of LLMs on emotion-controllable TTS tasks remains underexplored. To bridge this gap, we conduct a comparative analysis to evaluate the influence of LLM on model performance with different EmoVoice variants. Specifically, we keep the backbone architecture consistent, with the only variation being whether LLM parameters are used for initialization.

### v4

@@ 6.1.1. Objective evaluation results
The experimental results in Table 2 demonstrate that our proposed model, EmoVoice, achieves competitive performance across all evaluated metrics. Compared with existing emotional speech synthesis models, EmoVoice(1.5B) achieves the most balanced overall performance, with highest emotion similarity of 0.9118, and highest recall rate of 0.424, while maintaining low WER of 2.62 and high UTMOS score of 4.345. It rivals the powerful GPT-4o-mini-tts in content consistency and emotional expressiveness, while surpassing it in terms of speech quality. These results indicate that EmoVoice not only preserves linguistic accuracy and delivers high perceptual quality, but also excels in producing emotionally rich speech that faithfully represents the intended emotional directives. GPT-4o-audio serves as an upper bound reference rather than a comparable model.
@@ 6.1.2. Subjective evaluation results
Table 3. Comparison of Mean Opinion Score (MOS) for emotional expressiveness and instruction-following capabilities of different emotion-controllable TTS models.
@@
Given the inherently subjective and abstract nature of emotional expressiveness, objective metrics can only provide limited insight into model performance. Therefore, we conduct a subjective assessment to better align evaluations with human perception. Specifically, we randomly sample 10 utterances form each of six emotion categories from the test set, totaling 60 audio samples. These samples are evaluated by 30 distinct participants, each of whom rates 10 random groups of speech produced by six representative models. The evaluation focuses on two key aspects: the overall emotional expressiveness of the synthesized speech and its adherence to the given emotional instructions. Ratings are scored on a scale of 1 to 5, with increments of 0.5.
The results presented in Table 3 reveal that the MOS trends align closely with the objective emotion metrics, validating the consistency of our findings. Notably, the GPT-4o series models exhibit the strongest performance, with GPT-4o-audio achieving the highest MOS of 4.35, reflecting its ability to produce highly compelling and emotionally rich speech. Our EmoVoice models demonstrate competitive performance, surpassing both CosyVoice2 and PromptTTS by a significant margin. Specifically, the EmoVoice(1.5B) achieves a MOS of 3.507, closely approaching the performance of GPT-4o-mini-tts of 3.598 and outperforming CosyVoice2 of 2.138 and PromptTTS of 1.415, reaffirming the effectiveness of our approach in producing emotionally expressive and instruction-aligned speech.
@@ 6.1.3. Chinese model results
Table 4. Performance comparison of EmoVoice and other emotion-controllable TTS models on Secap test set.
@@ 6.2.1. Different output-side structures
Table 5. Performance comparison of EmoVoice variants with different output structures on EmoVoice-DB testset. Both Fine-tuned Emotional TTS and Pre-trained TTS are evaluated. WER of the Pre-trained TTS is also computed on test-en of Seed-TTS (Anastassiou et al., 2024).
@@
EmoVoice-$P$ Parallel Phoneme and Audio Tokens 3.06 0.9115 0.379 4.35 3.94/3.11 0.8704 0.298 4.40
@@
We explore optimizing the output-side structure of the EmoVoice model, comparing six different architectures, including standard single-output stream and parallel output streams, as shown in Figure 2. We investigate utilizing text and phonemes as auxiliary information to guide the generation of audio tokens. The baseline EmoVoice model outputs a pure audio token sequence. The EmoVoice-I model produces an interleaved sequence of text and audio tokens, mixed in a predefined ratio of 12:36, a structure commonly employed in streaming TTS models (Du et al., 2024b; Yang et al., 2024b). Inspired by CoT and MoT techniques, we experiment with first predicting text or phoneme tokens as an auxiliary intermediate thought process before generating the target audio token within a single output stream. These models are denoted as EmoVoice-ST(Serial Text) and EmoVoice-SP(Serial Phoneme), respectively. Besides, we explore using an additional prediction head in the output structure to produce text or phoneme tokens simultaneously. EmoVoice-PT(Parallel Text) and EmoVoice-PP(Parallel Phoneme) represent models that parallelly predict audio tokens alongside text tokens and phoneme tokens, respectively.
@@ 6.2.3. Scaling LLM size
Given the well-established scaling law in LLMs, we explore how scaling the parameter size of the LLM backbone affects the model’s performance on the emotional TTS task. Specifically, we compare two different sizes of 0.5B and 1.5B parameters, using the qwen2.5-0.5B and qwen2.5-1.5B models respectively.
As shown in Tabel 7, in the pre-training phase, we observe an obvious reduction of WER when scaling up the model. The 1.5B model achieves a WER of 3.83% with a relative reduction of 19.03% compared to the 4.73% WER of the 0.5B model, indicating enhanced content consistency. During the emotion fine-tuning phase, the recall rate rises from 0.395 to 0.424 and the WER reduces from 2.73% to 2.62%. Emotion similarity and UTMOS remain largely consistent. Additionally, subjective MOS score in Table 3 improves from 3.163 to 3.507. Overall, scaling up the LLM effectively enhances emotional expressiveness and content consistency. However, it incurs substantial additional computational costs and training time, with improvements constrained by the limited emotion dataset size.
@@ 6.2.4. LLM Initialization
While several TTS models adopt LLMs as their backbone and utilize their parameters for initialization, the specific impact of LLMs on emotion-controllable TTS tasks remains underexplored. To bridge this gap, we conduct a comparative analysis to evaluate the influence of LLM with different EmoVoice variants. Specifically, we keep the backbone architecture consistent, with the only variation being whether LLM parameters are used for initialization.
@@ 8. Conclusions
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No. U23B2018 and No. 62206171), Shanghai Municipal Science and Technology Major Project under Grant 2021SHZDZX0102 and Yangtze River Delta Science and Technology Innovation Community Joint Research Project (2024CSJGG01100).